Use any of the models( any of appendices 3.1 to 3.7) for representing fractions to find the sum of the following fractions:

a. 1/6 + 5/6

b. 1/5 + 3/5

c. 2/9 + 5/9

d. 1/8 + 2/8

e. 3/12 + 5/12


1
Expert's answer
2021-07-12T13:22:20-0400

a.


"\frac{1}{6} + \frac{5}{6}= \frac{1+5}{6}=\frac{6}{6}=1"



b.


"\frac{1}{5} + \frac{3}{5}= \frac{1+3}{5}=\frac{4}{5}"



c.


"\frac{2}{9} + \frac{5}{9}= \frac{2+5}{9}=\frac{7}{9}"



d.


"\frac{1}{8} + \frac{2}{8}= \frac{1+2}{8}=\frac{3}{8}"



e.


"\frac{3}{12} + \frac{5}{12}= \frac{3+5}{12}=\frac{8}{12}"